OpenStudy (anonymous):

lim x approaches pi/2+ cosx/1-sinx use L'Hospitals rule

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/2^+}\frac{ cosx }{ 1-sinx }\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so this gives us the indeterminate form 0/0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just don't understand y the answer equals - infinity

OpenStudy (anonymous):

so we take the derivative with respect to x of the numerator and then of the denominator

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i got \[\lim_{x \rightarrow \pi/2+} tanx\]

OpenStudy (anonymous):

i just don't understand y that equals \[-\]

OpenStudy (anonymous):

- infinity

OpenStudy (anonymous):

you mean the original limit? have you tried graphing it?

OpenStudy (anonymous):

i answer for this problem is - infinity but when u plug in pi/2+ for x in tanx don't u use the unit circle so tan pi/2 is undefined right

OpenStudy (anonymous):

y is it -infinity

OpenStudy (anonymous):

yes pi/2 is undefined because as this function approaches pi/2 from the right, it goes to negative infinity

OpenStudy (anonymous):